Scores on a test are approximately normally distributed with a mean of 70 and a standard deviation of 9. The teacher wants to give Aʹs to the top 10% of students, Bʹs to the next 25%, and Cʹs to the next 42%. What is the bottom cutoff for a C grade? Round your answer to the nearest whole number.
A) 77
B) 65
C) 63
D) 68

Answer by Boreal(15235)

Bottom cutoff for C is after top 77% or 23rd percentile.
That is at z=-0.74
-0.74=(x-mean)/sd
9*-0.74=x-70
63.34=x
C.